If there are three times as many red balls are there are yellow and half as many yellow balls as blue and there are 84 balls in total how many red vall are there
Answer by ewatrrr(24785)

 
Hi,
let x represent number of blue ⇒ x/2 yellow and 3x/2 red
x + x/2 + 3x/2 = 84
2x + x + 3x = 168
6x = 168
x = 28, number of blue. 14 yellow and 42 red
CHECKING our answer***
28 + 14 + 42 = 84
Wish You the Best in your Studies.